# Mitigating the Modality Gap: Few-Shot Out-of-Distribution Detection with Multi-modal Prototypes and Image Bias Estimation

### v1

@@ 5.1 Experimental Details
Datasets and benchmarks. Our experiments primarily use the ImageNet-1k [7] and ImageNet-100 [7] dataset as ID data. Aligning with standards from prior works [45, 61], we evaluate on four diverse OOD datasets: iNaturalist [59], SUN [63], Places [70], and Texture [6]. Additionally, we test on ImageNet-10 [7] and ImageNet-20 [7] for near OOD. Details are in the Appendix.
Baselines methods. We compare suPreMe with 19 baseline methods, including traditional OOD detection methods [21, 12, 41, 18, 38, 60, 56], VLM (CLIP)-based OOD methods [2, 45, 4, 61, 57, 46, 40, 48, 36, 24], and visual prompt tuning methods [73, 72].
Implementation details. All the images in the official training set are used to obtain ID image prototypes. For each class, 16 images are used for few-shot tuning. We use the image and text encoders of VIT-B/16 pre-trained by CLIP [53] for all the experiments. The parameters of CLIP are frozen. We use SGD to optimize other parameters, e.g., estimated image domain bias and learnable context, with a momentum of $0.9$. Training epochs, learning rate, batch size, and learnable context length are set to $50,0.002,32$, and $16$, respectively. We set $\alpha=0.005$ and $\beta=0.1$. Experiments are conducted on a single Nvidia V100 with 32 GB memory. All the reported results for suPreMe are the average of three trials.
@@ 5.2 Main Results
ImageNet-1k and ImageNet-100 as the ID dataset. In Tab. 3, we present a comprehensive analysis on ImageNet-1k (ID) across four OOD datasets (iNaturalist, SUN, Places, Texture) using CLIP (VIT-B/16). suPreMe demonstrates superior performance, achieving the lowest FPR95 and highest AUROC scores across most datasets. Notably, suPreMe achieves an average FPR95 of 20.28 and an average AUROC of 95.54. In Tab. 3, we observe consistent trends using ImageNet-100 as the ID dataset. suPreMe achieves the best average results, with FPR95 of 8.19 and AUROC of 98.08, showing substantial improvements over other methods. These results underscore our method’s robustness across varied OOD benchmark datasets, emphasizing its effectiveness in enhancing model reliability for real-world applications. However, we note that these OOD datasets are mostly semantically different from the ID data (ImageNet), making this problem easier. To evaluate suPreMe under more challenging conditions (near OOD), we follow previous work [45, 24] and conduct experiments on ImageNet-10 and ImageNet-20.
@@
FPR95 (%) $\downarrow$ AUROC (%) $\uparrow$ FPR95 (%) $\downarrow$ AUROC (%) $\uparrow$
[…]
@@
NegLabel 5.10 98.86 4.60 98.81
suPreMe 4.00 98.93 3.80 98.98
Table 4: Comparison on ImageNet-10 $\leftrightarrow$ ImageNet-20. We use CLIP-B/16 for all the experiments. Best in bold. “*” represents our reproduction. $\dagger$” refers to results from EOE [4].
@@
ImageNet-10 $\leftrightarrow$ ImageNet-20. In Tab. 4, we compare OOD detection performance tested with ImageNet-10 as the ID dataset and ImageNet-20 as the OOD dataset (ImageNet-10 $\rightarrow$ ImageNet-20), and vice versa (ImageNet-20 $\rightarrow$ ImageNet-10). suPreMe achieves the best performance, with the lowest FPR95 and highest AUROC scores in both scenarios. These results underscore its robustness and reliability in near OOD.
5.3 Ablation Study
To understand the effectiveness of suPreMe, we present the average results of an ablation study using CLIP (ViT-B/16) on ImageNet-1k $\rightarrow$ four OOD datasets. The detailed results are given in the Appendix.
@@ 5.3 Ablation Study
Effectiveness of the proposed modules (BPG and ITC) and scores ($S_{\textit{GMP}}$). In Tab. 6, we evaluate the effectiveness of BPG and ITC, as well as the proposed score $S_{\textit{GMP}}$, for enhancing OOD detection. We see that excluding BPG (-BPG) raises FPR95 to 27.51 and reduces AUROC to 93.71, highlighting BPG’s importance in reducing false positives. Similarly, removing ITC (-ITC) significantly degrades performance, increasing FPR95 to 28.59 and lowering AUROC to 93.38. Furthermore, when the model uses $S_{\textit{MCM}}$ instead of $S_{\textit{GMP}}$, the performance also declines, with FPR95 at 25.03 and AUROC at 94.80. Our full model, incorporating ITC, BPG, and $S_{\textit{GMP}}$, achieves the best overall performance. This demonstrates the combined efficacy of these modules and the proposed OOD score $S_{\textit{GMP}}$.
@@
suPreMe improves the ID performance. OOD detection acts as a preliminary step before models make predictions, aiming to perform a binary classification to distinguish between ID and OOD samples. However, it is also important to understand how OOD methods work with ID classification. In Tab. 6, we compare the Top-1 accuracy of different methods on the ImageNet-1K validation set. suPreMe achieves the highest accuracy, reaching 71.8%. Compared to other OOD detection methods, e.g., CLIPN (68.5%) and ID-like (68.3%), suPreMe shows a considerable gain, underscoring its effectiveness in maintaining high ID accuracy while also excelling in OOD detection.
Figure 3: Performance comparison on the length of prompts.
Impact of the length of prompts. In the main experiments, we follow previous works [46, 72, 73] and set the length of prompts to $L=16$. Here, we would like to validate that $16$ is the best choice for suPreMe. We see in Fig. 4 that among the choices from $1$ to $16$, $16$ obtains the lowest FPR95 and highest AUROC. Increasing the length of prompts improves the OOD detection performance.
Impact of base image size for constructing ID image prototypes. In $S_{\textit{GMP}}$, we use a set of base images to construct image prototypes $\{P_{i,c}\}_{c\in[C]}$. To understand how the size of base images affects performance, we conduct experiments varying its size. The results in Fig. 4 show that by increasing the number of images for each class, the performance of suPreMe is improved, as image prototypes are closer to the real distribution. FPR95 and AUROC reach the best when using all the images.
@@
Effectiveness of $S_{\textit{GMP}}$. To empirically show the effectiveness of $S_{\textit{GMP}}$, we visualize the empirical cumulative distribution of $S_{\textit{MCM}}$ and our $S_{\textit{GMP}}$ in Fig. 5. We use our best model trained on ImageNet-1k (ID), with iNaturalist being the OOD dataset. $S_{\textit{GMP}}$ clearly improves the gap between ID and OOD data in the empirical CDF figure. Moreover, the statistics of Kolmogorov–Smirnov test is enlarged from $0.7270$ ($S_{\textit{MCM}}$) to $0.8724$ ($S_{\textit{GMP}}$), indicating that using $S_{\textit{GMP}}$ has more separable scores compared to $S_{\textit{MCM}}$.
@@
(b) $S_{\textit{GMP}},\textit{KS}=0.8724$.
Figure 5: Comparison between $S_{\textit{MCM}}$ [45] and our $S_{\textit{GMP}}$ on ImageNet-1k (ID) to iNaturalist (OOD). The scores are multiplied by $100$ for better illustration. “KS” is the statistic from the Kolmogorov–Smirnov test. Higher KS statistic values indicate a greater difference between two distributions. Best viewed in color.
@@ 6 Conclusion
In this paper, we introduced suPreMe, a novel VLM-based few-shot tuning OOD detection framework designed to minimize the modality gap between text and image in CLIP. suPreMe incorporates two key modules: BPG, which enhances image-text fusion, and ITC, which minimizes the inter- and intra-modal distances. We also proposed the generalized multi-modal prototype OOD score, $S_{\textit{GMP}}$, using multi-modal prototypes and embeddings to improve robustness and reduce false positives. Experimental results show that suPreMe consistently outperforms existing VLM-based OOD detection methods.
Limitations. The efficiency of suPreMe partially relies on the introduction of multi-modal prototypes, which may not always be feasible in scenarios with strict data constraints. Exploring more advanced methods such as image generation or retrieval from a web-scale dataset could help alleviate the dependency on the ID data. Moreover, our suPreMe and previous VLM-based methods use CLIP as the base model. It would be interesting to explore the efficacy of using advanced VLMs, such as ImageBind [14], in OOD detection. Finally, in this work we have only explored the effectiveness of multi-modal positive/ID prototypes. Introducing negative/OOD prototypes may further enhance our framework’s versatility.
@@
In this Appendix, we present the details and complete results corresponding to Tab. 1 in Sec. 7.1, additional ablation studies in Sec. 7.2, benchmark dataset specifications in Sec. 7.3, and the proof of Theorem 1 in Sec. 8.
@@ 7.1 Empirical Evidence for RQ1
ImageNet-100 $\rightarrow$ iNaturalist
$S_{\textit{MCM}}$ [45] 18.13 96.77
$S_{\textit{MMP}}$ (Ours, Eq. 4) 14.76 97.36
ImageNet-100 $\rightarrow$ SUN
$S_{\textit{MCM}}$ [45] 36.45 94.54
$S_{\textit{MMP}}$ (Ours, Eq. 4) 30.28 92.51
ImageNet-100 $\rightarrow$ Places
$S_{\textit{MCM}}$ [45] 34.52 94.36
$S_{\textit{MMP}}$ (Ours, Eq. 4) 34.04 93.92
ImageNet-100 $\rightarrow$ Texture
$S_{\textit{MCM}}$ [45] 41.22 92.25
$S_{\textit{MMP}}$ (Ours, Eq. 4) 17.66 96.66
@@
14.64 96.76 22.92 95.01 33.05 91.86 42.55 90.00 28.29 93.40
18.75 96.15 24.47 94.43 33.36 91.80 42.96 89.92 29.89 93.07
20.00 96.02 22.35 95.65 28.19 93.82 29.57 93.82 25.03 94.80
12.73 97.27 23.86 94.80 33.65 91.67 39.80 91.10 27.51 93.71
15.78 96.79 24.56 94.23 33.35 91.55 40.66 90.93 28.59 93.38
8.27 98.29 19.40 95.84 26.69 93.56 26.77 94.45 20.28 95.54
Table 8: Effectiveness of different modules and the proposed OOD score $S_{\textit{GMP}}$. When $S_{\textit{GMP}}$ is not combined with BPG, it degenerates to $S_{\textit{MMP}}$.
@@
To empirically justify the effectiveness of multi-modal (image and text) prototypes, we present experiments using $S_{\textit{MCM}}$ and $S_{\textit{MMP}}$ on four OOD datasets (i.e., iNaturalist, SUN, Places, and Texture). We use CLIP-B/16 as the base model for $S_{\textit{MCM}}$ and $S_{\textit{MMP}}$. All the training images from ImageNet-100 are used for generating image prototypes.
Our method, $S_{\textit{MMP}}$, consistently demonstrates superior performance in terms of FPR95 across all four tasks, achieving the lowest values, with notable improvements on iNaturalist (14.76 vs. 18.13) and Texture (17.66 vs. 41.22), highlighting its robustness in reducing false positives. For AUROC, $S_{\textit{MMP}}$ achieves the highest score on iNaturalist (97.36) and Texture (96.66). However, in SUN and Places datasets, while $S_{\textit{MMP}}$ maintains competitive performance, it slightly underperforms $S_{\textit{MCM}}$ in AUROC (92.51 vs. 94.54 on SUN and 93.92 vs. 94.36 on Places).

### v2

@@ 5.1 Experimental Details
Datasets and benchmarks. Our experiments primarily use the ImageNet-1k [6] and ImageNet-100 [6] dataset as ID data. Aligning with standards from prior works [44, 59], we evaluate on four diverse OOD datasets: iNaturalist [57], SUN [61], Places [71], and Texture [5]. OpenOOD benchmark [68] is also used for evaluation on ImageNet-1k. We test on ImageNet-10 [6] and ImageNet-20 [6] for hard OOD. Details are in the Appendix.
Baselines methods. We compare suPreMe with several baseline methods [21, 11, 41, 18, 38, 58, 54, 44, 59, 24, 3, 55, 45, 40, 47, 2, 36, 66, 67, 74, 73].
Implementation details. For each class, 16 images are used for few-shot tuning and the calculation of ID image prototypes. We use the image and text encoders of VIT-B/16 pre-trained by CLIP [51] for all the experiments. To avoid overfitting, instead of using deep MLPs, we employ linear transformations without bias as the mappings ($f_{\textit{img-txt}}$ and $f_{\textit{txt-img}}$). The parameters of CLIP are frozen. We use SGD to optimize other parameters, e.g., estimated image domain bias and learnable context, with a momentum of $0.9$. Training epochs, learning rate, batch size, and learnable context length are set to $50,0.002,32$, and $16$, respectively. We set $\alpha=0.005$ and $\beta=0.1$. Experiments are conducted on a single Nvidia V100 with 32 GB memory. All the results for suPreMe are the average of three trials.
@@ 5.2 Main Results
ImageNet-1k and ImageNet-100 as the ID dataset. In Tab. 2, we present a comprehensive analysis of ImageNet-1k (ID) across four OOD datasets. suPreMe demonstrates superior performance, achieving the lowest FPR95 and highest AUROC scores across most datasets. Notably, suPreMe achieves an average FPR95 of 20.70 and an average AUROC of 95.38. With all the training data for calculating ID prototypes, suPreMe is further boosted with an average FPR95 of 20.28 and an average AUROC of 95.54. In Tab. 3, we observe consistent trends using ImageNet-100 as the ID dataset. suPreMe achieves the best average results, with FPR95 of 8.47 and AUROC of 98.08, showing substantial improvements over other methods.
ImageNet-1k on OpenOOD. The results on the OpenOOD benchmark are shown in Tab. 5. Compared to previous methods, suPreMe achieves the best performance in both near and far OOD scenarios. Specifically, suPreMe achieves an FPR95 of 71.44 and AUROC of 75.95 in the near OOD scenario, and an FPR95 of 20.92 and AUROC of 95.45 in the far OOD scenario.
ImageNet-10 $\leftrightarrow$ ImageNet-20. To evaluate suPreMe under challenging conditions where ID and OOD datasets share similar semantics, we follow previous work [44, 24] and conduct experiments on ImageNet-10 and ImageNet-20 as shown in Tab. 5. suPreMe achieves the best performance, with the lowest FPR95 and highest AUROC scores in both scenarios, demonstrating its robustness.
5.3 Ablation Study
We present an ablation study using CLIP (ViT-B/16) on ImageNet-1k $\rightarrow$ four OOD datasets. The computational complexity analysis (Tab. 14), compatibility with post-hoc methods (Tab. 12), impacts of shot count (Fig. 8), hyperparameter sensitivity (Figs. 10, 9 and 7), and the use of out-of-distribution prototypes (Tab. 12) are in the Appendix.
Methods NearOOD FarOOD
FPR95 $\downarrow$ AUROC $\uparrow$ FPR95 $\downarrow$ AUROC $\uparrow$
MCM 85.37 58.97 69.87 77.11
LoCoOp 87.95 64.97 59.88 83.23
suPreMe 71.44 75.95 20.92 95.45
Table 4: OOD Detection results with ID data of ImageNet-1k on the OpenOOD benchmark. Best in bold.
@@ 5.3 Ablation Study
FPR95 $\downarrow$ AUROC $\uparrow$ FPR95 $\downarrow$ AUROC $\uparrow$
Energy 10.30 97.94 16.40 97.37
CLIPN 7.80 98.07 13.67 97.47
MCM 5.00 98.71 17.40 98.87
@@
suPreMe 4.10 98.96 3.60 99.05
Table 5: Comparison on ImageNet-10 $\leftrightarrow$ ImageNet-20 with CLIP-B/16. Best in bold. “*” represents our reproduction.
@@
$\left\|\overrightarrow{\Delta}_{\text{gap}}\right\|$ ImageNet (ID) $\downarrow$ iNaturalist $\downarrow$ Places $\downarrow$ Texture $\downarrow$ SUN $\downarrow$
MCM 0.8679 0.9839 0.9101 0.9373 0.9252
suPreMe 0.6135 0.7159 0.8961 0.9273 0.9002
Table 8: Quantifying the modality gap with different methods. We use CLIP (ViT-B/16) as the base model.
@@
Effectiveness of BPG, ITC, and $\boldsymbol{S_{\textit{GMP}}}$. In Tab. 8, we evaluate the effectiveness of BPG and ITC, as well as the proposed score $S_{\textit{GMP}}$, for enhancing OOD detection. We see that excluding BPG (-BPG) raises FPR95 to 28.12 and reduces AUROC to 92.77, highlighting BPG’s importance in reducing false positives. Similarly, removing ITC (-ITC) significantly degrades performance, increasing FPR95 to 28.74 and lowering AUROC to 93.27. We also observe that the Gaussian bias term in ITC is important, as evidenced by the drop in performance (fourth row). Furthermore, when the model uses $S_{\textit{MCM}}$ instead of $S_{\textit{GMP}}$, the performance also declines, with FPR95 at 26.02 and AUROC at 94.79. Our full model, incorporating ITC, BPG, and $S_{\textit{GMP}}$, achieves the best performance. The detailed table is in Tab. 11 in the Appendix.
Using MLP for image-text consistency. $f_{\textit{img-txt}}$ and $f_{\textit{txt-img}}$ in ITC can sufficiently address the modality gap, we replace them with a two-layer MLP (Linear-ReLU-Linear) and present the experimental results in Tab. 8 (w. MLP). Results indicate that using MLP negatively impacts performance, which might be because, with limited training data, sophisticated modules such as MLP could overfit.
@@
suPreMe improves the ID performance. In Tab. 8, we compare the Top-1 accuracy of different methods on the ImageNet-1K validation set. Compared to other OOD detection methods, e.g., CLIPN (68.5%) and ID-like (68.3%), suPreMe shows a considerable gain with the highest accuracy, demonstrating its effectiveness in maintaining high ID accuracy while also excelling in OOD detection.
@@
Effectiveness of $\boldsymbol{S_{\textit{GMP}}}$. We visualize the empirical cumulative distribution of $S_{\textit{MCM}}$ and our $S_{\textit{GMP}}$ in Fig. 5 with the ID data of ImageNet-1k and the OOD data of iNaturalist using the same model (suPreMe). $S_{\textit{GMP}}$ clearly improves the gap between ID and OOD data in the empirical CDF figure. Furthermore, the Kolmogorov–Smirnov test statistic is enlarged from 0.7270 ($S_{\textit{MCM}}$) to 0.8555 ($S_{\textit{GMP}}$), indicating that using $S_{\textit{GMP}}$ yields more separable scores than $S_{\textit{MCM}}$.
@@
(b) $S_{\textit{GMP}},\textit{KS}=0.8555$.
Figure 4: Comparison of (a) $S_{\textit{MCM}}$ [44] and (b) our $S_{\textit{GMP}}$ on ImageNet-1k (ID) to iNaturalist (OOD). The scores are multiplied by $100$ for better illustration. KS is the Kolmogorov–Smirnov statistic. Higher KS values indicate a greater difference between the distributions. Best viewed in color.
Figure 5: The t-SNE visualization of embeddings from two arbitrary classes in ImageNet-1k (Val) using suPreMe.
Reduced modality gap (t-SNE visualization and quantitative assessment of modality gap). We choose two arbitrary classes from ImageNet-1k (Val) and use t-SNE to visualize embeddings in Fig. 5. It is clear that ITC reduces the modality gap, as the distances between mapped image embeddings $I^{\prime}$ and text prototypes $P_{\textit{txt}}$ are smaller than that between $I$ and $P_{\textit{txt}}$. Similarly, we observe that image embeddings $I$ are located close to image prototypes $P_{\textit{img}}$. We also quantitatively assess the modality gap following Liang et al. [39] as shown in Tabs. 8 and 15, which shows that suPreMe achieves smaller modality gaps compared with previous methods.
@@ 6 Conclusion
We have introduced suPreMe, a novel VLM-based few-shot tuning OOD detection method. suPreMe incorporates two key modules: BPG, which enhances image-text fusion, and ITC, which minimizes the inter- and intra-modal distances. We also proposed the generalized multi-modal prototype OOD score, $S_{\textit{GMP}}$, using multi-modal prototypes and embeddings to improve robustness and reduce false positives. Experiments show that suPreMe consistently outperforms existing VLM-based OOD detection methods.
Limitations. In some scenarios, the ID data may not be available. Exploring methods such as image generation or retrieval from a web-scale dataset could help alleviate the dependency on the ID data. Moreover, exploring more advanced distribution models rather than the Gaussian distribution we use may further improve the performance.
Acknowledgement
This work was supported by the Natural Sciences and Engineering Research Council of Canada (NSERC)-CSE Research Community project entitled “An End-to-End Approach to Safe and Secure AI Systems”.
Researchers funded through the NSERC-CSE Research Communities Grants do not represent the Communications Security Establishment Canada or the Government of Canada. Any research, opinions, or positions they produce as part of this initiative do not represent the official views of the Government of Canada.
@@ Acknowledgement
(a) ImageNet $\rightarrow$ iNaturalist (OOD)
(b) ImageNet $\rightarrow$ Texture (OOD)
(c) ImageNet $\rightarrow$ Places (OOD)
(d) ImageNet $\rightarrow$ SUN (OOD)
Figure 6: Examples where OOD images from four OOD datasets are located closer to ID text prototypes compared to ID images (ImageNet-1k) generated with CLIP (ViT-B/16).
In this Appendix, we present the details and complete results corresponding to Tab. 1 in Sec. 8.1, full theoretical motivation in Sec. 8.2, additional ablation studies in Sec. 9.2, benchmark dataset specifications in Sec. 9.3, and the proof of Theorem 1 in Sec. 10.
7 OOD Examples
In Figs. 2 and 6, we visualize the embedding space using t-SNE projection across four different OOD datasets (iNaturalist, Texture, Places, and SUN) generated with CLIP (ViT-B/16). A pattern emerges: OOD images (red crosses) are consistently positioned closer to ID text prototypes than their corresponding ID images across all four scenarios. For instance, in Fig. 6(a), OOD images (red cross) are positioned closer to the “sulphur butterfly” text prototype (purple circle) than the corresponding ID images (blue pentagon). This counterintuitive positioning is systematically observed across diverse semantic categories—from biological entities (sulphur butterfly, starfish) to textures (honeycomb, cherimoya) and environmental scenes (jay/cougar, rapeseed/hay).
This phenomenon reveals a fundamental limitation in current embedding spaces: positional proximity does not guarantee semantic similarity. The t-SNE visualization demonstrates that geometric distance in the learned representation space can be misleading, as semantically unrelated OOD samples may appear closer to class prototypes than genuine in-distribution examples. This misalignment between spatial proximity and semantic similarity undermines the reliability of distance-based OOD detection methods. It motivates our design of multimodal prototypes that better capture the true semantic boundaries of in-distribution data.
8 RQ1
8.1 More Empirical Evidence for RQ1
Methods FPR95 (%) $\downarrow$ AUROC (%) $\uparrow$
ImageNet-100 $\rightarrow$ 4 datasets (Average)
Text prototypes 32.58 94.48
Image prototypes 31.09 94.74
Text + Image prototypes 24.18 95.79
ImageNet-100 $\rightarrow$ iNaturalist
Text prototypes 18.13 96.77
Image prototypes 26.06 93.09
Text + Image prototypes 14.76 97.36
ImageNet-100 $\rightarrow$ SUN
Text prototypes 36.45 94.54
Image prototypes 39.25 93.17
Text + Image prototypes 30.28 92.51
ImageNet-100 $\rightarrow$ Places
Text prototypes 34.52 94.36
Image prototypes 40.59 96.43
Text + Image prototypes 34.04 93.92
ImageNet-100 $\rightarrow$ Texture
Text prototypes 41.22 92.25
Image prototypes 18.46 96.27
Text + Image prototypes 17.66 96.66
To empirically justify the effectiveness of multi-modal (image and text) prototypes, we present experiments using $S_{\textit{MCM}}$ and $S_{\textit{MMP}}$ on four OOD datasets (i.e., iNaturalist, SUN, Places, and Texture) as shown in Tab. 9. We use CLIP-B/16 as the base model for $S_{\textit{MCM}}$ and $S_{\textit{MMP}}$. All the training images from ImageNet-100 are used for generating image prototypes. We also include $S_{\textit{MCM}}$ with image prototypes as a comparison baseline.
Our method, $S_{\textit{MMP}}$, consistently demonstrates superior performance in terms of FPR95 across all four tasks, achieving the lowest values, with notable improvements on iNaturalist (14.76 vs. 18.13) and Texture (17.66 vs. 41.22), highlighting its robustness in reducing false positives. For AUROC, $S_{\textit{MMP}}$ achieves the highest score on iNaturalist (97.36) and Texture (96.66). However, in the SUN and Places datasets, while $S_{\textit{MMP}}$ maintains competitive performance, it slightly underperforms $S_{\textit{MCM}}$ in AUROC (92.51 vs. 94.54 on SUN and 93.92 vs. 94.36 on Places).
Similarly, we observe that $S_{\textit{MMP}}$ outperforms $S_{\textit{MCM}}$ with image prototypes on most of the metrics, while $S_{\textit{MCM}}$ with image prototypes has slightly better performance compared to $S_{\textit{MCM}}$ (with text prototypes). This shows that the modality gap negatively impacts the performance.
8.2 Theoretical Evidence for RQ1
To further understand why image prototypes improve performance, we present a theoretical analysis in Theorem 1.
This greater score separation enhances the separation between ID and OOD data, improving detection performance, under general assumptions used by previous works [12, 39]. And it further improves performance. The proof is provided in Sec. 10.
@@ 9 Experiments
LoCoOp 85.60 60.73
suPreMe 64.97 75.32
Table 10: OOD Detection results with ID data of ImageNet-1k on the OpenOOD benchmark (CSID). Best in Bold.
@@
8.56 98.48 22.36 95.38 33.60 92.55 39.57 92.75 26.02 94.79
20.86 95.06 22.04 94.92 30.22 91.64 39.38 89.46 28.12 92.77
16.10 96.50 23.81 94.53 32.63 91.94 42.41 90.11 28.74 93.27
✓(w/o Bias) 21.55 97.79 16.37 98.00 26.25 92.24 34.25 91.56 24.61 94.90
5.91 98.73 15.85 96.08 26.18 93.19 34.88 91.93 20.70 95.38
w. ResNet-50 21.93 96.22 37.73 91.00 43.27 87.41 28.14 93.51 32.77 92.04
w. ResNet-101 6.13 98.58 20.61 95.03 33.16 90.94 41.31 90.70 25.30 93.81
Table 11: Effectiveness of different modules, the proposed OOD score $S_{\textit{GMP}}$, and different backbones. When $S_{\textit{GMP}}$ is not combined with BPG, it degenerates to $S_{\textit{MCM}}$.
@@
Methods ResNet-50 ResNet-101
FPR95 $\downarrow$ AUROC $\uparrow$ FPR95 $\downarrow$ AUROC $\uparrow$
NegPrompt 57.19 85.70 50.35 88.04
LoCoOp 34.45 91.95 - -
suPreMe 30.55 92.90 25.30 93.81
Table 13: Performance comparison using different base models.
@@
9.1 Results on OpenOOD (CSID)
We evaluate the performance of suPreMe on the OpenOOD benchmark (CSID) [21] using the ImageNet-1k dataset as the ID data in Tab. 10. The results show that suPreMe outperforms LoCoOp, achieving the lowest FPR95 and highest AUROC, demonstrating its effectiveness in OOD detection across different benchmarks.